OpenStudy (anonymous):

Solve for x. x+2/3 + 2x-4/4=3 4 9 correct? 4/9 9/2

OpenStudy (anonymous):

\[x+\frac{ 2 }{ 3 }+\frac{ 2x-4 }{ 4 }=3\] Like that?

OpenStudy (anonymous):

Your wrong then

OpenStudy (anonymous):